Operations Report – February 7th

Crew 221 Operations Report 07Feb2022

SOL: 5

Name of person filing report: Luis D. Monge

Non-nominal systems:

Notes on non-nominal systems:

Generator: off (at 8pm) waiting for SOC to reach 80% to be turned on

Hours run: 11:00

From what time last night: 8:00pm

To what time this morning: 7:00am

List any additional daytime hours when the generator was run : 0

Solar— SOC % (Before generator is run at night) 84%

Diesel Reading – 98%

Station Propane Reading – 60%

Ethanol Free Gasoline – gallons

Water (loft tank): 35 gallons

Water Meter: 148 units

Water (static tank) – 320 gals

Static to Loft Pump used – yes

Water in GreenHab – 232.5 gals

Water in ScienceDome: 0 gallons

Toilet tank emptied: no

Deimos rover used: no


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR


Beginning charge:

Ending charge:

Currently charging: yes

Spirit rover used: yes

Hours: (before EVA) : 129.5

Beginning charge: (Before EVA) 100%

Ending charge: (On return from EVA, before recharging) 63%

Currently charging: yes

Opportunity rover used: no

Hours: n/a

Beginning charge: n/a

Ending charge: n/a

Currently charging: yes

Curiosity rover used: yes

Hours: 135.7

Beginning charge: 100%

Ending charge: 45%

Currently charging: yes

Notes on rovers:

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) not used

Reason for use: not used

Oil Added? not used

ATV Fuel Used: Gals not used

# Hours the ATVs were Used today: not used

Notes on ATVs:

HabCar used and why, where? yes, for mail run by Atila

CrewCar used and why, where? not used

General notes and comments: not used

Summary of internet: operational

Summary of suits and radios: operational, lose cable on suit 7 fixed

Summary of Hab operations: nominal.

Summary of GreenHab operations: operational, black corn experiment continues watered with 4L. watering of the plants carried out as per procedure

Summary of ScienceDome operations: 2 incubators running, microbiology experiments going on and cultures are being grown in agar plates.

Summary of RAM operations: robot assembly is ongoing and expected to continue.

Summary of any observatory issues: nominal

Summary of health and safety issues: crew member had epistaxis, treated inmediately by crew doctor, HSO report submitted.

Questions, concerns and requests to Mission Support: every time the oven is turned on there is a heavy gas smell.

Copyright © The Mars Society. All rights reserved. | Main Site